# From Bias to Balance: Exploring and Mitigating Spatial Bias in LVLMs

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) have achieved remarkable success across a wide range of multimodal tasks, yet their robustness to spatial variations remains insufficiently understood. In this work, we present a systematic study of the spatial bias of LVLMs, focusing on how models respond when identical key visual information is placed at different locations within an image. Through a carefully designed probing dataset, we demonstrate that current LVLMs often produce inconsistent outputs under such spatial shifts, revealing a fundamental limitation in their spatial-semantic understanding. Further analysis shows that this phenomenon originates not from the vision encoder, which reliably perceives and interprets visual content across positions, but from the unbalanced design of position embeddings in the language model component. In particular, the widely adopted position embedding strategies, such as RoPE, introduce imbalance during cross-modal interaction, leading image tokens at different positions to exert unequal influence on semantic understanding. To mitigate this issue, we introduce Balanced Position Assignment (BaPA), a simple yet effective mechanism that assigns identical position embeddings to all image tokens, promoting a more balanced integration of visual information. Extensive experiments show that BaPA enhances the spatial robustness of LVLMs without retraining and further boosts their performance across diverse multimodal benchmarks when combined with lightweight fine-tuning. Further analysis of information flow reveals that BaPA yields balanced attention, enabling more holistic visual understanding.  [^1] Our data and code are available at https://github.com/AAAndy-Zhu/BaPA.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) have achieved remarkable success across a wide range of multimodal tasks, including visual question answering (Alayrac et al., 2022; Xu et al., 2025), image captioning (Li et al., 2023; Lu et al., 2025), and open-ended reasoning (Zhu et al., 2025b; Wang et al., 2025). By combining powerful vision encoders with large language models (LLMs), these systems are able to integrate information from both modalities and perform complex reasoning. Despite these advances, LVLM still exhibits fundamental limitations when it comes to spatially robust semantic understanding of visual content (Imam et al., 2025; Li et al., 2025).
@@
To uncover the root cause of this vulnerability, we first investigate whether the vision encoder is responsible for the observed spatial bias. Our eraser search (Li et al., 2016; De Cao et al., 2020) experiments on its perceptual ability confirm that it consistently perceives visual features of key content, regardless of the spatial position of the key image. Building on this, we then analyze whether the encoder’s semantic understanding is affected by spatial position. The high and stable similarity between text embeddings and visual features across different locations demonstrates that semantic encoding is also robust to spatial variation. These observations rule out the vision encoder as the source of spatial bias, suggesting that the issue originates in the LLM portion of the LVLM, where the visual features are processed for multimodal reasoning.
Based on the above observations, we hypothesize that spatial bias stems from the imbalance in cross-modal interactions, primarily introduced by position embeddings within the LLM backbone of LVLMs. Current models typically adopt RoPE (Su et al., 2024) or related schemes, which modulate attention scores between tokens based on their relative distances in the sequence. While this design has proven highly effective in unimodal text modeling, it becomes problematic in multimodal contexts. Specifically, all image tokens should contribute equally during cross-modal fusion with text tokens. The sequential distance bias inherent in RoPE disrupts this equitable interaction, leading to distorted cross-modal interaction and ultimately undermining the spatial robustness of LVLMs.
To address this, we propose a simple yet effective Balanced Position Assignment (BaPA) mechanism which assigns identical positional embeddings to all image tokens. This modification explicitly promotes a more balanced and thorough integration of visual cues in cross-modal interactions. Empirical results on our probing dataset show that applying BaPA without retraining leads to more balanced performance and remarkably higher accuracy. We further adapt BaPA for broader downstream tasks. Results on five benchmark datasets show that BaPA can enhance LVLMs across diverse multimodal benchmarks, demonstrating the effectiveness and generalizability of BaPA.
@@
To summarize, our contributions are threefold. First, we provide a systematic investigation of spatial robustness in LVLMs’s semantic understanding, empirically demonstrating through novel probes that LVLMs’ predictions vary significantly with the spatial location of identical visual content. Second, we pinpoint the root cause of this bias to the LLM’s position-sensitive cross-modal interactions, rather than to deficiencies in the vision encoder. Third, we introduce Balanced Position Assignment (BaPA), a lightweight and generalizable method that improves spatial robustness without sacrificing downstream performance, validated through both probing tasks and five multimodal benchmarks.
@@ 2 Related Work
Large Vision-Language Models. LVLMs combine both visual and textual inputs, providing a more comprehensive understanding of visual spatial relationships, objects, and scenes (Bordes et al., 2024). Existing LVLMs typically comprise a visual encoder (Dosovitskiy et al., 2021; Radford et al., 2021), a projector (Alayrac et al., 2022), and a pre-trained LLM (Touvron et al., 2023). Through pre-training with image-text pairs and fine-tuning with preference or instruction, current LVLMs, like LLaVA (Liu et al., 2023a) and Qwen2.5-VL (Bai et al., 2025), have been successful in dialogue (Zhu et al., 2025a), question answering (Zhu et al., 2023), and complex reasoning (Zhu et al., 2025b). Nonetheless, LVLMs still exhibit numerous biases (Ruggeri et al., 2023; Wang et al., 2024a; Zhang et al., 2025b), which diminish the trustworthiness of their response. In this paper, we systematically investigate the under-explored issue of spatial bias of LVLMs and introduce a mitigation strategy from the perspective of positional balance.
Position Encoding in Transformers. Since Transformers (Vaswani et al., 2017) lack a natural understanding of sequence order, numerous studies have proposed various position encoding methods. Early works adopt sinusoidal absolute position encoding (Vaswani et al., 2017) and learnable embeddings (Dosovitskiy et al., 2021), while later approaches focus on learnable and relative position encodings (Shaw et al., 2018; Huang et al., 2020; He et al., 2021; Ke et al., 2021), which offer more flexibility in capturing temporal dependencies. Among them, Rotary Position Embedding (RoPE) (Su et al., 2024) encodes relative positions through rotation matrices and has shown strong effectiveness in LLMs and LVLMs. Meanwhile, work on extensions to RoPE is emerging, such as LongRoPE (Ding et al., 2024), Unified RoPE (Wu et al., 2025), and CARoPE (Veisi et al., 2025). Despite this, current work find that the impact of RoPE leads to position bias in LLMs (Zhang et al., 2024; Chen et al., 2023), resulting in a phenomenon known as “lost-in-middle” (Liu et al., 2023b). Recently, Xing et al. (2024) investigate how RoPE affects object hallucination in LVLMs and propose a novel position alignment method to mitigate the long-term decay in RoPE. In contrast, we thoroughly explore the spatial bias grounded in the fundamental semantic understanding task and observe that the underlying issue lies not in long-term decay but in unbalanced position assignment of LVLMs. Although our approach also involves modifying the position encoding, it is motivated by a different rationale, and it is empirically more effective than Xing et al. (2024) on multiple benchmarks and even achieves balanced results without retraining.
3 Probing Task for Spatial Robustness
@@
In this section, we introduce the proposed probing task designed to systematically examine the spatial robustness of LVLM’s semantic understanding ability. Specifically, we assess whether model predictions remain consistent when key information appears in different regions of an image. We first introduce the design of our probing task, and then present experimental results to reveal the vulnerabilities of LVLMs on spatial-semantic understanding.
@@ 3.1 Task Design
To systematically evaluate the spatial robustness of LVLMs, we construct a probe dataset based on image–text matching. Specifically, we randomly sample 10,000 image–caption pairs $(I_{m},C_{m})$ from the LAION dataset (Schuhmann et al., 2022). As shown in Figure 6, for each key image $I_{m}$, we first randomly retrieve 8 distractor images from LAION. We then arrange $I_{m}$ and the 8 distractors in a $3\times 3$ grid to form a composite image $\bm{I}_{m,0}$, which is presented to the LVLM together with the caption $C_{m}$. The model is asked a yes-or-no question $Q_{m}$ to determine whether any sub-image within the composite matches the given caption $C_{m}$. To probe sensitivity to spatial variation, we further construct augmented composites $\{\bm{I}_{m,1},\bm{I}_{m,2},\dots,\bm{I}_{m,8}\}$, where the original image $I_{m}$ is placed at different grid locations. In each case, the same question $Q_{m}$ is posed to the model. The final dataset contains 90,000 samples $\{\bm{I}_{m,n},C_{m},Q_{m}|m\in\{0,1,\dots,9,999\},n\in\{0,1,\dots,8\}\}$ in total. By comparing outputs across these variants, we can directly assess whether LVLMs yield consistent predictions in response to positional changes of key information, while all other visual and textual factors remain unchanged. More details are available in Appendix B.
Figure 1: Results on our probe dataset, where each cell reports the model’s accuracy when the key image $I_{m}$ is placed at a specific grid location $n$ in the $3\times 3$ composite image $\bm{I}_{m,n}$.
3.2 Results and Findings.
@@
We evaluate five representative LVLMs on our probing dataset, including Qwen2.5-VL-7B(Bai et al., 2025), Gemma3-12B (Team et al., 2025), LLaVA-v1.6-Mistral-7B (LLaVA-v1.6) (Liu et al., 2024), and Llama3-LLaVA-NeXT-8B (LLaVA-NeXT). All models are evaluated in a zero-shot setting, without any task-specific fine-tuning.
@@
Main Results. As shown in Figure 1, all LVLMs exhibit sensitivity to the spatial variation of the key image. LLaVA-v1.6, and particularly LLaVA-NeXT exhibit strong sensitivity to spatial variations, with large fluctuations in accuracy across positions. Among all models, Qwen2.5-VL achieves the most consistent performance, likely benefiting from its improved MRoPE (Bai et al., 2025), although slight spatial bias can still be observed. Moreover, we observe that the performance is not well correlated with token distance, suggesting that spatial bias might not stem from the long-term decay property of RoPE (Su et al., 2024).
@@
Impact of Model Scale. We further examine the correlation between model scale and spatial bias by evaluating the Qwen2.5-VL model series on our probe dataset. As shown in Table 1, Qwen2.5-VL-7B exhibits greater accuracy fluctuations across different grid positions ($\Delta=1.74$), reflecting higher spatial sensitivity and stronger bias. With increasing model size, Qwen2.5-VL-32B generally achieves lower variance, which is consistent with previous observations (Zhu et al., 2024) that larger models tend to make more consistent decisions. Interestingly, the average accuracy of Qwen2.5-VL-32B declines slightly compared to Qwen2.5-VL-7B, the reason can be that larger models might be more susceptible to overfitting or inefficiencies in processing specific data types.
@@ 4 Analyzing the Origin of Spatial Bias
To pinpoint the origin of the spatial bias in LVLMs, this section systematically examines the role of the vision encoder. Our analysis focuses on two critical aspects of the encoder: its low-level visual perception and high-level visual understanding capabilities, aiming to ascertain if the bias stems from this component.
4.1 Is Vision Encoder Perception Spatially Robust?
Task Design. Inspired by the work on model interpretability (Si et al., 2024; Zhang et al., 2025a), we design a set of experiments using eraser search (Li et al., 2016; De Cao et al., 2020) based on our probe dataset, where each region of the input image are occluded sequentially and we observe the resulting changes in the model’s response. This allows us to examine how the model’s behavior is affected by different spatial locations of key image $I_{m}$ in the composite input $\bm{I}_{m,n}$. In specific, for each position $n$, we first randomly select 20 key images, with a total of 180 samples. As shown in Figure 8(a), each composite image is divided into 400 non-overlapping regions, and we mask each region one at a time by perturbing its pixels to the background color (white in our experiment). Then, we calculate the difference between the logits of generated token based on the original composite image and the masked version for each LVLM and take it as the importance score of the region with respect to the response. Finally, we aggregate these importance scores into a heatmap to visualize how information from different positions within the input image influences the model’s decision.
(a) Results on perception ability across each Position $n$ of key image. Darker regions indicate higher importance scores, where logits change more significantly before and after masking. More results of Qwen2.5-VL and LLaVA-v1.6 are available in Figure 8(a).
Results and Findings. Figure 3(a) presents the results of our masking experiments on Gemma3 and LLaVA-NeXT. In both models, the vision encoder is consistently able to identify the critical regions corresponding to the key image, irrespective of its location in the composite input. In addition, the overall patterns remain semantically aligned with the key image and do not shift when its position changes, demonstrating that the vision encoder’s perception is robust to spatial variation. These results rule it out as the source of the spatial bias. More results are presented in Figure 8(a).

### v2

# Beyond the Vision Encoder: Identifying and Mitigating Spatial Bias in Large Vision-Language Models

@@ Abstract
Large Vision-Language Models (LVLMs) have achieved remarkable success across a wide range of multimodal tasks, yet their robustness to spatial variations remains insufficiently understood. In this work, we conduct a systematic study of the spatial bias of LVLMs, examining how models respond when identical key visual information is placed at different locations within an image. Through controlled probing experiments, we observe that current LVLMs often produce inconsistent outputs under such spatial shifts, revealing a clear spatial bias in their semantic understanding. Further analysis indicates that this bias does not stem from the vision encoder, but rather from a mismatch in attention mechanisms between the vision encoder and the large language model, which disrupts the global information flow. Motivated by this insight, we propose Adaptive Global Context Injection (AGCI), a lightweight mechanism that dynamically injects shared global visual context into each image token. AGCI works without architectural modifications, mitigating spatial bias by enhancing the semantic accessibility of image tokens while preserving the model’s intrinsic capabilities. Extensive experiments demonstrate that AGCI not only enhances the spatial robustness of LVLMs, but also achieves strong performance on various downstream tasks and hallucination benchmarks.
Machine Learning, ICML
@@ 1 Introduction
Large Vision-Language Models (LVLMs) have achieved remarkable success across a wide range of multimodal tasks, including visual question answering (Alayrac et al., 2022; Xu et al., 2025), image captioning (Li et al., 2023a; Lu et al., 2025), and open-ended reasoning (Zhu et al., 2025b; Wang et al., 2025a). By combining powerful vision encoders with large language models (LLMs), these systems are able to integrate information from both modalities and perform complex reasoning. Despite these advances, LVLM still exhibits fundamental limitations when it comes to spatially robust semantic understanding of visual content (Imam et al., 2025; Li et al., 2025b; Qi et al., 2025).
@@
To uncover the root cause of this vulnerability, we examine LVLMs from two stages of visual information processing: perception and semantic understanding (Li et al., 2025a). Our eraser search (Li et al., 2016; De Cao et al., 2020) experiments on LVLM’s perceptual ability confirm that it consistently perceives visual features of key content, regardless of the spatial position of the key image. Building on this, we further analyze whether the vision encoder’s semantic understanding is robust to spatial position. The high and stable similarity between text embeddings and visual features across different locations suggest that semantic encoding is also robust to spatial variation. These observations rule out the perceptual ability and vision encoder as the source of spatial bias, suggesting that the issue originates in the LLM portion of the LVLM, where the visual features are processed for multimodal reasoning.
Based on the above observations, we hypothesize that spatial bias stems from the mechanism mismatch between the bidirectional self-attention used in vision encoders (typically ViT-based (Dosovitskiy et al., 2021) in LVLM) and the causal attention employed in LLMs. While vision encoders allow image tokens to exchange contextual information globally and symmetrically, the causal constraint in LLMs fundamentally disrupts this interaction pattern: earlier tokens cannot access subsequent contextual information as originally designed, thereby breaking the symmetrical global information flow. Consequently, during cross‑modal reasoning, the contribution of each image token becomes overly dependent on its position—manifesting as the observed spatial bias.
To address this, we propose Adaptive Global Context Injection (AGCI), a lightweight and effective mechanism that explicitly reinforces global visual semantics at the LLM stage. In specific, AGCI injects the global visual context into each image token based on semantic similarity, strengthening tokens with less information while preserving those that are already semantically aligned. This design restores stable access to global context of image during cross-modal reasoning without modifying the underlying model architecture. We adapt AGCI for broader downstream tasks. Results on six benchmark datasets show that AGCI can enhance LVLMs not only on general VQA, but also on hallucination‑ and OCR-oriented tasks, demonstrating the effectiveness and generalizability of AGCI. And the visualization of information flow from image token to text token further emphasizes the importance of incorporating global image context.
@@
To summarize, our contributions are threefold. First, we provide a systematic investigation of spatial robustness in LVLMs’s semantic understanding, empirically demonstrating through novel probes that LVLMs’ predictions vary significantly with the spatial location of identical visual content. Second, we pinpoint the root cause of this bias to the loss of global image context caused by the causal attention mechanism in LLMs, rather than to deficiencies in the vision encoder. Third, we introduce Adaptive Global Context Injection (AGCI), a simple yet effective method that improves spatial robustness without sacrificing downstream performance, validated through both probing tasks and six multimodal benchmarks [^1] Our data and code are available at https://github.com/AAAndy-Zhu/AGCI.
Figure 1: Results on our probe dataset, where each cell reports the model’s accuracy when the key image $I_{m}$ is placed at a specific grid location $n$ in the $3\times 3$ composite image $\boldsymbol{I}_{m,n}$.
@@ 2 Related Work
Large Vision-Language Models. LVLMs combine both visual and textual inputs, providing a more comprehensive understanding of visual spatial relationships, objects, and scenes (Bordes et al., 2024). Existing LVLMs typically comprise a visual encoder (Dosovitskiy et al., 2021; Radford et al., 2021), a projector (Alayrac et al., 2022), and a pre-trained LLM (Touvron et al., 2023). Through pre-training with image-text pairs and fine-tuning with preference or instruction, current LVLMs, like LLaVA (Liu et al., 2023) and Qwen2.5-VL (Bai et al., 2025b), have been successful in dialogue (Zhu et al., 2025a), question answering (Zhu et al., 2023), and complex reasoning (Zhu et al., 2025b). Nonetheless, LVLMs still exhibit numerous biases (Ruggeri et al., 2023; Wang et al., 2024a; Zhang et al., 2025b), which diminish the trustworthiness of their response. In this paper, we systematically investigate the under-explored issue of spatial bias of LVLMs and introduce a mitigation strategy from the perspective of global image context.
Position Bias. Position bias has been widely studied in LLMs The pioneering work reveals that LLMs typically suffer from “the lost in the middle” phenomenon when processing multiple documents (Liu et al., 2024b). Similar biases have been observed in LLM-as-a-judge scenarios, where models favor earlier responses regardless of their quality (Zheng et al., 2023; Shi et al., 2025). To address this issue, recent works have proposed various mitigation strategies for LLMs (Hsieh et al., 2024; Yu et al., 2025; Wang et al., 2025b). As LVLMs inherit the autoregressive LLM backbone, these position biases naturally extend to multimodal settings. However, existing studies on position bias in LVLMs remain limited. In multi-image reasoning, prior work demonstrates that LVLM predictions are highly sensitive to image ordering and introduces a training-free mitigation method based on interpolating causal and bidirectional inter-image attention (Tian et al., 2025). Beyond this, several studies on multi-modal in-context learning also indirectly reveal the effects of position in LVLMs, showing that images placed at the beginning contribute minimally to model predictions (Chen et al., 2024; Baldassini et al., 2024). In the single-image setting, recent efforts have begun to explore spatial bias within images, often in the context of object hallucination. Xing et al. (2024) first investigate how RoPE affects object hallucination in LVLMs when the object placed at different spatial positions in the image and propose a novel position alignment method to mitigate the long-term decay in RoPE. Zhu et al. (2025c) then argue that such explanations lack generality and propose attention calibration mechanisms to rebalance spatial attention. In contrast, our work provides a principled analysis grounded in the fundamental semantic understanding task and introduce an effective mitigation strategy for spatial bias in single-image scenario.
3 Probing Spatial Robustness in Semantic Understanding: A Controlled Study
@@
This section presents a controlled probing task designed to systematically evaluate the spatial robustness of LVLMs’ semantic understanding ability. Specifically, we assess whether model predictions remain consistent when key information appears in different regions of an image. We first introduce the design of our probing task, and then present experimental results to reveal the vulnerabilities of LVLMs on spatial-semantic understanding.
@@ 3.1 Task Design
To systematically evaluate the spatial robustness of LVLMs, we construct a probe dataset based on image–text matching. Specifically, we randomly sample 10,000 image–caption pairs $(I_{m},C_{m})$ from the LAION dataset (Schuhmann et al., 2022). As shown in Figure 6, for each key image $I_{m}$, we first randomly retrieve 8 distractor images from LAION. We then arrange $I_{m}$ and the 8 distractors in a $3\times 3$ grid to form a composite image $\boldsymbol{I}_{m,0}$, which is presented to the LVLM together with the caption $C_{m}$. The model is asked a yes-or-no question $Q_{m}$ to determine whether any sub-image within the composite matches the given caption $C_{m}$. To probe sensitivity to spatial variation, we further construct augmented composites $\{\boldsymbol{I}_{m,1},\boldsymbol{I}_{m,2},\dots,\boldsymbol{I}_{m,8}\}$, where the original image $I_{m}$ is placed at different grid locations. In each case, the same question $Q_{m}$ is posed to the model. The final dataset contains 90,000 samples $\{\boldsymbol{I}_{m,n},C_{m},Q_{m}|m\in\{0,1,\dots,9,999\},n\in\{0,1,\dots,8\}\}$ in total, with each image $\boldsymbol{I}_{m,n}$ featuring a resolution of $840\times 840$. By comparing outputs across these variants, we can directly assess whether LVLMs yield consistent predictions in response to positional changes of key information, while all other visual and textual factors remain unchanged. More details are available in Appendix A.
3.2 Results and Findings
@@
We evaluate six representative LVLMs on our probing dataset, including Qwen2.5-VL-7B (Bai et al., 2025b), Qwen3-VL-8B-Instruct (Bai et al., 2025a), Gemma3n-E4B-it (Team et al., 2025), LLaVA-v1.6-Mistral-7B (LLaVA-v1.6) (Liu et al., 2024a), and InternVL3-8B. All models are evaluated in a zero-shot setting, without any fine-tuning.
@@
Main Results. As shown in Figure 1, all LVLMs exhibit sensitivity to the spatial variation of the key image, particularly LLaVA-v1.6, with large fluctuations in accuracy across positions. Among all models, Qwen2.5-VL achieves the most consistent performance, likely benefiting from its improved MRoPE (Bai et al., 2025b), although slight spatial bias can still be observed. Interestly, despite sharing a similar architecture, Qwen3-VL shows a more pronounced bias, possibly attributed to the enhanced Interleaved MRoPE failing to leverage its advantages in single-image scenarios. Moreover, we observe that the performance is not well correlated with token distance, suggesting that spatial bias might not stem from the long-term decay property of RoPE (Su et al., 2024). This observation is further supported by experiments with varying resolutions of $\boldsymbol{I}_{m,n}$, where changing the number of image tokens does not lead to notable differences in the bias pattern, reinforcing that this issue is unlikely to be caused by RoPE. The results are presented in Appendix B.
(a) Results on perception ability across each Position $n$ of key image. Darker regions indicate higher importance scores, where logits change more significantly before and after masking. More results are available in Figure 11(a).
@@
Impact of Model Scale. We further examine the correlation between model scale and spatial bias by evaluating the Qwen2.5-VL model series on our probe dataset. As shown in Table 1, Qwen2.5-VL-7B exhibits greater accuracy fluctuations across different grid positions ($\Delta=1.74$), reflecting higher spatial sensitivity and stronger bias. With increasing model size, Qwen2.5-VL-32B and Qwen2.5-VL-72B generally achieves lower variance, which is consistent with previous observations (Zhu et al., 2024) that larger models tend to make more consistent decisions. However, the average accuracy tends to decrease, the reason can be that larger models might be more susceptible to overfitting or inefficiencies in processing specific data types.
@@ 4 Analyzing the Origin of Spatial Bias
To pinpoint the origin of the spatial bias in LVLMs, we following previous work (Li et al., 2025a) and analyze LVLMs by decomposing their visual information processing into two stages: perception and semantic understanding.
Figure 3: Understanding abilities of vision encoders across each Position $n$. Each cell reports the similarity between visual features and text embeddings input to the LLMs when the key image $I_{m}$ is placed at a specific grid location $n$ in the $3\times 3$ composite image $\boldsymbol{I}_{m,n}$.
4.1 Is LVLM Perception Spatially Robust?
Task Design. Inspired by the work on model interpretability (Si et al., 2024; Zhang et al., 2025a), we design a set of experiments using eraser search (Li et al., 2016; De Cao et al., 2020) based on our probe dataset, where each region of the input image are occluded sequentially and we observe the resulting changes in the model’s response. This allows us to examine how the model’s behavior is affected by different spatial locations of key image $I_{m}$ in the composite input $\boldsymbol{I}_{m,n}$. In specific, for each position $n$, we first randomly select 20 key images, with a total of 180 samples. As shown in Figure 7(a), each composite image is divided into 400 non-overlapping regions, and we mask each region one at a time by perturbing its pixels to the background color (white in our experiment). Then, we calculate the difference between the logits of generated token based on the original composite image and the masked version for each LVLM and take it as the importance score of the region with respect to the response. Finally, we aggregate these importance scores into a heatmap to visualize how information from different positions within the input image influences the model’s decision.
Results and Findings. Figure 3(a) presents the results of our masking experiments on Gemma3n and InternVL3. Both models are consistently able to identify the critical regions corresponding to the key image, irrespective of its location in the composite input. In addition, the overall patterns remain semantically aligned with the key image and do not shift when its position changes, demonstrating that the LVLM’s perception is robust to spatial variation. These results rule it out as the source of the spatial bias. More results are presented in Figure 11(a).